OpenStudy (anonymous):

how do you choose the point on the terminal side of -45 degrees?

hero (hero):

I don't know what point you're talking about but 360 - 45 = 315

OpenStudy (anonymous):

I think the question is asking what are the coordinates on the unit circle of -45 degrees..I don't really understand the question

hero (hero):

Well, you should be more specific next time. You should ask what are the `coordinates` of a point on a `unit circle` where \(\theta\) equals \(-45^\circ\).

hero (hero):

And the answer to that is to simply find \((x,y) = (\sin(315^\circ\)), \(\cos(315^\circ))\)

OpenStudy (anonymous):

Oh okay thank you so much!!!!